OpenStudy (anonymous):

(8 multiplied by sq. root of 2) - (sq. root of 2) pls show step by step

myininaya (myininaya):

\[8\sqrt{2}-\sqrt{2}=8\sqrt{2}-1\sqrt{2}=\sqrt{2}(8-1)=(8-1)\sqrt{2}=7\sqrt{2}\]

OpenStudy (anonymous):

nice use of the "distributive property" combined with the "commutative law"

OpenStudy (anonymous):

although... you could have written \[8\sqrt{2}-\sqrt{2}=8\sqrt{2}-1\sqrt{2}=(8-1)\sqrt{2}=7\sqrt{2}\]